# Refining Pre-Trained Motion Models

### 01 Jan 2024 v1

@@ Abstract
Given the difficulty of manually annotating motion in video, the current best motion estimation methods are trained with synthetic data, and therefore struggle somewhat due to a train/test gap. Self-supervised methods hold the promise of training directly on real video, but typically perform worse. These include methods trained with warp error (i.e., color constancy) combined with smoothness terms, and methods that encourage cycle-consistency in the estimates (i.e., tracking backwards should yield the opposite trajectory as tracking forwards). In this work, we take on the challenge of improving state-of-the-art supervised models with self-supervised training. We find that when the initialization is supervised weights, most existing self-supervision techniques actually make performance worse instead of better, which suggests that the benefit of seeing the new data is overshadowed by the noise in the training signal. Focusing on obtaining a “clean” training signal from real-world unlabelled video, we propose to separate label-making and training into two distinct stages. In the first stage, we use the pre-trained model to estimate motion in a video, and then select the subset of motion estimates which we can verify with cycle-consistency. This produces a sparse but accurate pseudo-labelling of the video. In the second stage, we fine-tune the model to reproduce these outputs, while also applying augmentations on the input. We complement this boot-strapping method with simple techniques that densify and re-balance the pseudo-labels, ensuring that we do not merely train on “easy” tracks. We show that our method yields reliable gains over fully-supervised methods in real videos, for both short-term (flow-based) and long-range (multi-frame) pixel tracking.
@@ I Introduction
Self-supervised (or “unsupervised”) models hold the promise of training directly on real video, yet currently lag behind the supervised counterparts. State-of-the-art self-supervised motion methods typically rely on warp error (i.e., color constancy) combined with smoothness terms [6, 7, 8], and cycle-consistency (i.e., tracking backwards should yield the opposite trajectory as tracking forwards) [9, 10, 11, 12]. In current experimental setups, fully-supervised versions are typically positioned in a league ahead, reflecting in some way an unfair advantage. This divide between supervised and unsupervised methods is purely academic; if synthetic supervision is available and helpful, it makes sense to use it.
In this work, we take on the challenge of improving state-of-the-art supervised models with self-supervised training in an unabelled test domain. Unlike typical self-supervised or semi-supervised motion estimation approaches, which initialize a model from scratch, our setup is to take an arbitrary pre-trained model off the shelf, and attempt to improve it for a given test domain, without access to any additional data. We argue that this setup reflects a very practical use-case: a user notices some errors in the model’s output, and is willing to invest some GPU cycles into tuning up the results—provided that the tuning is effective.
@@
Focusing on the problem of obtaining a “clean” training signal from real-world unlabelled video, we propose to separate label-making and training into two distinct stages. In the first stage, we use the pre-trained model to estimate motion in a video, and then select the subset of motion estimates which we can verify with cycle-consistency. This produces a sparse but accurate pseudo-labelling of the video. In the second stage, we fine-tune the model to reproduce these outputs, while also applying augmentations on the input. We complement this self-training method with simple techniques that densify and re-balance the pseudo-labels, ensuring that we do not merely train on "easy" tracks.
@@ II Related Work
Learning motion from color and smoothness. Soon after the introduction of deep-learned flow methods [1], there has been interest in doing the learning self-supervised [6], converting optical flow’s classic assumptions [18], color constancy, and motion smoothness, into supervision objectives. Color constancy implies that when flow is accurate, corresponding pixels in consecutive frames should have the same color. By warping images to align with estimated flow, reducing per-pixel differences can improve flow accuracy, particularly when computed across multiple scales  [19]. The smoothness assumption helps to resolve ambiguities and occlusions which can be simply implemented by penalizing spatial gradients on the output flow. Minute details in this setup can affect performance greatly [20]. Instead of relying on color and smoothness, which can be understood as indirectly encouraging correct motion, we compute pseudo-labels for unlabelled data, providing direct regression targets.
@@
Training from pseudo labels. The technique of training on self-generated estimates (i.e., pseudo labels) was first proposed by Lee et al. [29]. This core idea is often paired with methods to filter the training targets to "confident" ones, where confidence may be approximated with the help of ensemble methods [30, 31]. Our work computes confidence with the help of domain knowledge: if a trajectory is within some cycle-consistency margin and color constancy, we consider it a pseudo-label. Similar to Chen et al. [31], we assume that it is possible to train on a small number of confident samples and generalize to other regions of the domain.
@@ III-B First Stage: Pseudo-label Generation
2-Frame Cycle-Consistency Filter. For optical flow, we compute consistency by warping the backward flow map into alignment with the forward flow map and then measure per-pixel disagreements with a sum (since opposite flows should cancel out) [32]. Denoting the forward flow as $w$ and the aligned backward flow as $\hat{w}$, we use:
$||w+\hat{w}||_{2}^{2}<\alpha(||w||_{2}^{2}+||\hat{w}||_{2}^{2}+\beta),$ (1)
@@
where a threshold of $1$ means that the trajectories may only drift a maximum of 1 pixel apart. Tracks that satisfy Eqn. 3 are considered cycle-consistent.
For both 2-frame and 8-frame tracking, the threshold values ($\tau$, $\alpha$, and $\beta$) depend on meta information such as the spatial resolution of the video. We measure the effect of these thresholds in our ablation study, and find that there is a Goldilocks zone: with thresholds set too small, yielding strict filter, there are too few good tracks for the pre-trained model to finetune; with thresholds set too large, yielding a loose filter, there are too many noisy tracks which corrupt the pre-trained weights.
@@ III-C Second Stage: Motion Refinement
In the second stage, we try to improve the pre-trained motion model by fine-tuning it on the collected pseudo labels, challenging the model to reproduce those motion estimates under more difficult settings. It may seem strange to fine-tune the model using its own estimates, but there are several reasons this may be helpful. First, the model is iterative (in the case of both RAFT and PIPs), and we are fine-tuning all iterations with the model’s final iteration, making the early iterations more confident to jump toward the answer. Second, taking gradient steps on these pseudo-labels may lead the model to “forget” behavior which is inconsistent with the pseudo-labels, which is a form of domain adaptation. Third, we apply augmentations while fine-tuning, which widens the domain where the model should produce these cycle-consistent outputs. We use the default augmentations from RAFT and PIPs, which include color jittering, cropping, resizing, and random occlusions.
@@
Figure 1: Refining a pre-trained motion model. We apply a pre-trained motion model on an unlabelled video, yielding a dense set of tracks. We filter this down to a sparse set of cycle-consistent tracks used as a pseudo-labelled dataset to finetune the model and improve its performance on the video.
@@ IV-B Evaluation Metrics
CroHD [15]. For CroHD, we used five different evaluation metrics to comprehensively evaluate the model performance. ATE-Vis and ATE-Occ evaluate the average track error in pixels each time with an $8$-frames video, for targets that stay fully-visible and ones that undergo occlusion, respectively. The other three metrics MTE, $\delta$, and Survival Rate, focus on longer-frame tracking. MTE denotes the median track error in pixels; $\delta$ is an accuracy metric measuring the ratio of points in our predictions that are within a threshold distance of ground truth, averaged over thresholds $(1,2,4,8,16)$. Survival measures the average time that a track manages to stay “on target” (within a threshold distance 16), measured as a ratio of video length [5].
@@ IV-C Baselines
TABLE I: Evaluation in CroHD. Ours clearly surpasses baselines in most of the metrics and improves ATE of pretrained PIPs[14] by $10\%$. Existing self-supervision methods make performance worse.
@@ IV-D Results
CroHD [15]. We use PIPs [14] pretrained on FlyingThings [2] dataset and run CroHD inference during sampling with a spatial resolution of $(768,1280)$. The 8-frame cycle consistency filter (Eqn. 3) and multiple-video fine-tuning (Sec. III-C) are adopted. We choose $2.5$ for the threshold $\tau$ and $3000$ for finetuning iterations $\kappa$.
@@
Figure 2: Performance change in $\delta$ for each of the video compared with the baseline pretrained PIPs model. Positive value denotes improvement (i.e., accuracy increasing)
Figure 3: Percent change in EPE for each video compared with the pretrained RAFT [26] model. Negative value denotes improvement (i.e., error decreasing).
@@
Results and comparisons are provided in Table I. We compare with state-of-the-art methods supervised out-of-domain [26, 34, 14], and also methods that perform self-supervised finetuning on a pre-trained model like ours. Our method clearly surpasses all of the baselines with better scores on most of the metrics. Notably, compared with pretrained PIPs where our optimization begins, we reduce ATE-Vis and ATE-Occ from $4.57$ and $7.71$ to $4.02$ and $6.79$, improving by around $10\%$ with our self-supervised refinement procedure.
For baselines, we also try several self-supervised techniques to finetune the pretrained model as ours, namely the PIPs+Color and PIPS+Feat shown in Table I. The idea of them is to use a self-supervised loss measuring the difference between the color or the image features at each predicted track position. We could observe that both these two techniques are not effective, demonstrating even worse scores on many evaluation metrics than the pretrained model. This further implies the effectiveness of our proposed procedure in finetuning a pretrained motion model.
Horse30 [16] We use PIPs [14] pretrained on FlyingThings [2] dataset and run Horse30 inference during sampling with a spatial resolution of $(256,448)$. The 8-frame cycle consistency filter (Eqn. 3) and multiple-video fine-tuning (Sec. III-C) are adopted. We choose $3$ for the threshold $\tau$ and $6000$ for finetuning iterations $\kappa$.
Our results on this dataset are shown in Table III. Compared to the pretrained PIPs [14], our method demonstrates a marked improvement in ATE and MTE. Specifically, we observed a decrease in ATE from $12.31$ to $10.81$ and in MTE from $4.40$ to $3.97$ with our self-supervised motion refinement technique. For reference, we also include the performance of pretrained RAFT [26] model, which registers an ATE of $12.46$ and MTE of $4.03$. We also include the PIPs+Color baseline implemented in the same way as in CroHD to showcase that the traditional motion self-supervision technique fails on pretrained model. Concretely, it achieves an ATE of $12.15$ and MTE of $4.40$, indicating that it barely helps to improve model performance. These results showcase the effectiveness of our method.
TABLE III: Evaluation in Horse30. Ours clearly surpasses baselines in all of the metrics and improves ATE of pretrained PIPs[14] by more than $10\%$. Existing self-supervision methods make performance worse.
@@
Compared with the CroHD results in Table I where we improve the pretrained model by $10\%$, results on Tap-Vid-DAVIS do not show as much improvement. However, we note that each Tap-Vid-DAVIS video is only $5$ seconds long on average (not much space to extract pseudo-labels), and ground truth is only available for a small number of points (less than 30 per video).
TABLE IV: Evaluation in MPI-Sintel. A clear advantage is observed for our method, improving pretrained RAFT by around $10\%$. Existing self-supervised methods do not help.
@@
MPI-Sintel. For MPI-Sintel, we use RAFT [26] pretrained on FlyingThings datasets and run Sintel inference during sampling with a spatial resolution of (436, 1024). We use the 2-frame cycle consistency filter (Eqn. 1) augmented by color consistency (Eqn. 2) and fine-tune on each video individually. We use the threshold hyperparameters $\alpha=0.005$, $\beta=0.25$, and $\gamma=0.1$. We choose $100$ for finetuning iterations $\kappa$ for each video. We evaluate performance change on each individual scene, and also measure the average across videos.
Quantitative comparisons are provided in Table IV, and qualitative comparisons are in Fig. 3, and Fig. 4. We compare with state-of-the-art methods supervised out-of-domain, self-supervised methods, and self-supervised techniques to finetune the pretrained model. We surpass all of the baselines, with better EPE on both clean and final splits of MPI-Sintel. Notably, compared with the pre-trained RAFT model, we reduce the EPE for both clean and final from $1.46$ and $2.72$ to $1.32$ and $2.46$, demonstrating around $10\%$ improvement. Since evaluation variance is quite huge on Sintel, we visualize the percentage EPE change for each of the videos in Figure 3. We can clearly observe that our self-supervised refinement improves the model performance on all of the scenes except ‘ambush_6’ in the clean split. On the other scenes, we observe obvious and significant improvements, even $30\%$ on scenes like ‘ambush_7’. We visualize the optical flow results in Fig. 4, and highlight the noticeably different regions in red boxes. We observe that our proposed method helps refine the optical flow output by completing missing objects, and improving smoothness in large regions.
As baselines to our optimization strategy, we try using color constancy as a self-supervised loss to finetune the pretrained model, optionally combined with a smoothness regularizer or edge-aware smoothness regularizer. We list these in Table IV as RAFT+Color, RAFT+Col.+Smooth and RAFT+Col.+Edge-Sm. In practice, we found that none of these techniques help, and EPE simply rises as training progresses.

### 17 Feb 2024 v2

@@ Abstract
Given the difficulty of manually annotating motion in video, the current best motion estimation methods are trained with synthetic data, and therefore struggle somewhat due to a train/test gap. Self-supervised methods hold the promise of training directly on real video, but typically perform worse. These include methods trained with warp error (i.e., color constancy) combined with smoothness terms, and methods that encourage cycle-consistency in the estimates (i.e., tracking backwards should yield the opposite trajectory as tracking forwards). In this work, we take on the challenge of improving state-of-the-art supervised models with self-supervised training. We find that when the initialization is supervised weights, most existing self-supervision techniques actually make performance worse instead of better, which suggests that the benefit of seeing the new data is overshadowed by the noise in the training signal. Focusing on obtaining a “clean” training signal from real-world unlabelled video, we propose to separate label-making and training into two distinct stages. In the first stage, we use the pre-trained model to estimate motion in a video, and then select the subset of motion estimates which we can verify with cycle-consistency. This produces a sparse but accurate pseudo-labelling of the video. In the second stage, we fine-tune the model to reproduce these outputs, while also applying augmentations on the input. We complement this boot-strapping method with simple techniques that densify and re-balance the pseudo-labels, ensuring that we do not merely train on “easy” tracks. We show that our method yields reliable gains over fully-supervised methods in real videos, for both short-term (flow-based) and long-range (multi-frame) pixel tracking. Our code can be found here: https://github.com/AlexSunNik/refining-motion-code.
@@ I Introduction
Self-supervised (or “unsupervised”) models hold the promise of training directly on real video, yet lag behind their supervised counterparts. State-of-the-art self-supervised motion methods typically rely on warp error (i.e., color constancy) combined with smoothness terms [6, 7, 8], and cycle-consistency (i.e., tracking backwards should yield the opposite trajectory as tracking forwards) [9, 10, 11, 12]. In current experimental setups, fully-supervised versions are typically positioned in a league ahead, reflecting in some way an unfair advantage. This divide between supervised and unsupervised methods is purely academic; if synthetic supervision is available and helpful, it makes sense to use it.
In this work, we improve state-of-the-art supervised models with self-supervised training in an unabelled test domain. Unlike typical self- or semi-supervised motion estimation approaches which are trained from scratch, we refine an off-the-shelf pre-trained model for a given test domain without requiring any additional data.
We argue that this setup reflects a very practical use-case: a user notices some errors in the model’s output, and is willing to invest some GPU cycles into tuning up the results—provided that the tuning is effective.
@@
In order to obtain a “clean” training signal from real-world unlabelled video, we propose to separate labeling and training into two distinct stages. In the first stage, we use the pre-trained model to estimate motion in a video, and then select the subset of motion estimates which we can verify with cycle-consistency. This produces a sparse but accurate pseudo-labeling of the video. In the second stage, we fine-tune the model to reproduce these outputs, while also applying augmentations on the input. We complement this self-training method with simple techniques that densify and re-balance the pseudo-labels, ensuring that we do not merely train on “easy” tracks.
@@ II Related Work
Learning motion from color and smoothness. Soon after the introduction of deep-learned flow methods [1], there has been interest in learning optical flows in self-supervised manner [6], converting classic flow assumptions [18], color constancy, and motion smoothness, into supervision objectives. Color constancy implies that when flow is accurate, corresponding pixels in consecutive frames should have the same color. By warping images to align with estimated flow, reducing per-pixel differences can improve flow accuracy, particularly when computed across multiple scales  [19]. The smoothness assumption helps to resolve ambiguities and occlusions which can be simply implemented by penalizing spatial gradients on the output flow. Minute details in this setup can affect performance greatly [20]. Instead of relying on color and smoothness, which can be understood as indirectly encouraging correct motion, we compute pseudo-labels for unlabelled data, providing direct regression targets.
@@
Training from pseudo labels. The technique of training on self-generated estimates (i.e., pseudo labels) was first proposed by Lee et al. [28]. This core idea is often paired with methods to filter the training targets to “confident” ones, where confidence may be approximated with the help of ensemble methods [29, 30]. Our work computes confidence with the help of domain knowledge: if a trajectory is within some cycle-consistency margin and color constancy, we consider it a pseudo-label. Similar to Chen et al. [30], we assume that it is possible to train on a small number of confident samples and generalize to other regions of the domain.
@@ III-B First Stage: Pseudo-label Generation
2-Frame Cycle-Consistency Filter. For optical flow, we compute consistency by warping the backward flow map into alignment with the forward flow map and then measure per-pixel disagreements with a sum (since opposite flows should cancel out) [31]. Denoting the forward flow as $w$ and the aligned backward flow as $\hat{w}$, $\forall$ pixel locations $i,j$, we check:
$||w_{ij}+\hat{w}_{ij}||_{2}^{2}<\alpha(||w_{ij}||_{2}^{2}+||\hat{w}_{ij}||_{2}^{2}+\beta),$ (1)
@@
where a threshold of $\tau=1$ means that the trajectories may only drift a maximum of 1 pixel apart. Tracks that satisfy Eqn. 3 are considered cycle-consistent.
For both 2-frame and 8-frame tracking, the threshold values ($\tau$, $\alpha$, and $\beta$) depend on meta information such as the spatial resolution of the video. We measure the effect of these thresholds in our ablation study, and find that there is a Goldilocks zone: with thresholds set too small, yielding a strict filter, there are too few good tracks for the pre-trained model to finetune; with thresholds set too large, yielding a loose filter, there are too many noisy tracks which corrupt the pre-trained weights.
@@ III-C Second Stage: Motion Refinement
In the second stage, we try to improve the pre-trained motion model by fine-tuning it on the batched and augmented pseudo labels, challenging the model to reproduce those motion estimates under more difficult settings. It may seem strange to fine-tune the model using its own estimates, but there are several reasons this may be helpful. First, the model is iterative (in the case of both RAFT and PIPs), and we are fine-tuning all iterations with the model’s final iteration, making the early iterations more confident to jump toward the answer. Second, taking gradient steps on these pseudo-labels may lead the model to “forget” behavior which is inconsistent with the pseudo-labels, which is a form of domain adaptation. Third, we apply augmentations while fine-tuning, which widens the domain where the model should produce these cycle-consistent outputs. We use the default augmentations from RAFT and PIPs, which include color jittering, cropping, resizing, and random occlusions.
@@
Figure 1: Refining a pre-trained motion model. We apply a pre-trained motion model on an unlabelled video, yielding a dense set of tracks. We filter this down to a sparse set of cycle-consistent tracks, creating pseudo-labels. We then train on the pseudo-labelled data using augmentations, to improve the model.
@@ IV-B Evaluation Metrics
CroHD [15]. For CroHD, we used five different evaluation metrics to comprehensively evaluate the model performance. ATE-Vis and ATE-Occ evaluate the average track error in pixels each time with an $8$-frames video, for targets that stay fully-visible and ones that undergo occlusion, respectively. The other three metrics MTE, $\delta$, and Survival Rate, focus on longer-frame tracking. MTE denotes the median track error in pixels; $\delta$ is an accuracy metric measuring the ratio of points in our predictions that are within a threshold distance of ground truth, averaged over thresholds $(1,2,4,8,16)$. Survival measures the average time that a track manages to stay “on target” (within a threshold distance 16) without falling off the trajectory and switching to other object or points, measured as a ratio of video length [5].
@@ IV-C Baselines
TABLE I: Evaluation in CroHD. Ours clearly surpasses baselines in most of the metrics and improves ATE of pre-trained PIPs[14] by $10\%$. Existing self-supervision methods make performance worse.
@@ IV-D Results
CroHD [15]. We use PIPs [14] pre-trained on FlyingThings [2] dataset and run CroHD inference during sampling with a spatial resolution of $(768,1280)$. The 8-frame cycle consistency filter (Eqn. 3) and multiple-video fine-tuning (Sec. III-C) are adopted. We choose $2.5$ for the threshold $\tau$ and $3000$ for finetuning iterations $\kappa$.
@@
Figure 2: Evaluation in Tap-Vid-DAVIS. Performance change in $\delta$ for each of the video compared with the baseline pre-trained PIPs model. Positive value denotes improvement (i.e., accuracy increasing)
Figure 3: Evaluation in MPI-Sintel. Percent change in EPE for each video compared with the pre-trained RAFT [13] model. Negative value denotes improvement (i.e., error decreasing).
@@
Results and comparisons are provided in Table I. We compare with state-of-the-art methods supervised out-of-domain [13, 33, 14], and also methods that perform self-supervised finetuning on a pre-trained model like ours. Our method clearly surpasses all of the baselines with better scores on most of the metrics. Notably, compared with pre-trained PIPs where our optimization begins, we reduce ATE-Vis and ATE-Occ from $4.57$ and $7.71$ to $4.02$ and $6.79$, improving by around $10\%$ with our self-supervised refinement procedure.
For baselines, we also try several self-supervised techniques to finetune the pre-trained model as ours, namely the PIPs+Color and PIPS+Feat shown in Table I. The idea of them is to use a self-supervised loss measuring the difference between the color or the image features at each predicted track position. We could observe that both these two techniques are not effective, demonstrating even worse scores on many evaluation metrics than the pre-trained model. This further implies the effectiveness of our proposed procedure in finetuning a pre-trained motion model.
Horse30 [16] We use PIPs [14] pre-trained on FlyingThings [2] dataset and run Horse30 inference during sampling with a spatial resolution of $(256,448)$. The 8-frame cycle consistency filter (Eqn. 3) and multiple-video fine-tuning (Sec. III-C) are adopted. We choose $3$ for the threshold $\tau$ and $6000$ for finetuning iterations $\kappa$.
Our results on this dataset are shown in Table III. Compared to the pre-trained PIPs [14], our method demonstrates a marked improvement in ATE and MTE. Specifically, we observed a decrease in ATE from $12.31$ to $10.81$ and in MTE from $4.40$ to $3.97$ with our self-supervised motion refinement technique. For reference, we also include the performance of pre-trained RAFT [13] model, which registers an ATE of $12.46$ and MTE of $4.03$. We also include the PIPs+Color baseline implemented in the same way as in CroHD to showcase that the traditional motion self-supervision technique fails on pre-trained model. Concretely, it achieves an ATE of $12.15$ and MTE of $4.40$, indicating that it barely helps to improve model performance. These results showcase the effectiveness of our method.
TABLE III: Evaluation in Horse30. Ours clearly surpasses baselines in all of the metrics and improves ATE of pre-trained PIPs[14] by more than $10\%$. Existing self-supervision methods make performance worse.
@@
Compared with the CroHD results in Table I where we improve the pre-trained model by $10\%$, results on Tap-Vid-DAVIS do not show as much improvement. However, we note that each Tap-Vid-DAVIS video is only $5$ seconds long on average (not much space to extract pseudo-labels), and ground truth is only available for a small number of points (less than 30 per video).
TABLE IV: Evaluation in MPI-Sintel. A clear advantage is observed for our method, improving pre-trained RAFT by around $10\%$. Existing self-supervised methods do not help.
@@
MPI-Sintel. For MPI-Sintel, we use RAFT [13] pre-trained on FlyingThings datasets and run Sintel inference during sampling with a spatial resolution of (436, 1024). We use the 2-frame cycle consistency filter (Eqn. 1) augmented by color consistency (Eqn. 2) and fine-tune on each video individually. We use the threshold hyperparameters $\alpha=0.005$, $\beta=0.25$, and $\gamma=0.1$. We choose $100$ for finetuning iterations $\kappa$ for each video. We evaluate performance change on each individual scene, and also measure the average across videos.
Quantitative comparisons are provided in Table IV and Fig. 3, and qualitative comparisons are in Fig. 4. We compare with state-of-the-art methods supervised out-of-domain, self-supervised methods, and self-supervised techniques to finetune the pre-trained model. We surpass all of the baselines, with better EPE on both clean and final splits of MPI-Sintel. Notably, compared with the pre-trained RAFT model, we reduce the EPE for both clean and final from $1.46$ and $2.72$ to $1.32$ and $2.46$, demonstrating around $10\%$ improvement. Since evaluation variance is quite huge on Sintel, we visualize the percentage EPE change for each of the videos in Figure 3. We can clearly observe that our self-supervised refinement improves the model performance on all of the scenes except ‘ambush_6’ in the clean split. On the other scenes, we observe obvious and significant improvements, even $30\%$ on scenes like ‘ambush_7’. We visualize the optical flow results in Fig. 4, and highlight the noticeably different regions in red boxes. We observe that our proposed method helps refine the optical flow output by completing missing objects, and improving smoothness in large regions.
As baselines to our optimization strategy, we try using color constancy as a self-supervised loss to finetune the pre-trained model, optionally combined with a smoothness regularizer or edge-aware smoothness regularizer. We list these in Table IV as RAFT+Color, RAFT+Col.+Smooth and RAFT+Col.+Edge-Sm. In practice, we found that none of these techniques help, and EPE simply rises as training progresses.